OpenStudy (anonymous):

Given that cos(theta) = 1/(square root7) and sin(theta)<0, determine the value of tan(theta)?

OpenStudy (anonymous):

instead of theta let me use x given cosx = 1/sqrt(7) and sinx<0 we need tanx

OpenStudy (anonymous):

secx=1/cosx =sqrt(7) (tanx)^2=(secx)^2-1=7-1=6